# Integrable Stochastic Processes Associated with the $D_{2}$ Algebra

### v1

@@ Abstract
We introduce novel integrable stochastic processes associated with the $D_{2}$ quantum group, which can be decomposed into two XXX spin chains (or two symmetric simple exclusion processes). We establish the integrability of the model under three types of boundary conditions (periodic, twisted, and open boundaries), and present its exact solution, including the spectrum, eigenstates, and some observables.
@@ 1 Introduction
In this work, we construct a novel integrable stochastic process associated with the $D_{2}$ spin chain [14, 15], which represents a prototypical quantum integrable system not based on an $A$-type Lie algebra, and has many applications in the high energy, topological and mathematical physics. This model incorporates four types of particles distinguished by their “charge” and “color”. Owing to the isomorphism $D_{2}\cong SU(2)\oplus SU(2)$, the system can be decomposed into two independent $SU(2)$ XXX spin chains. We demonstrate the integrability of the model under periodic, twisted, and open boundary conditions and derive its exact solution in the form of TTQQ relation, Bethe state and associated Bethe ansatz equations. The model introduced in this paper differs from existing multi-component stochastic processes with integrable boundary conditions, thereby enriching the family of exactly solvable stochastic systems.
The paper is structured as follows: In Section 2, Section 3, and Section 4, we study the stochastic process with periodic, twisted, and open boundary conditions, respectively. For each case, we demonstrate the integrability of the model and derive the corresponding exact solutions. The paper concludes with a summary and outlook in Section 5.
2 Integrable stochastic process with period boundary conditions
@@ 2.1 Transition matrix
When two particles of different colors occupy adjacent sites, they exchange positions with a symmetric transition rate, analogous to the dynamics in SSEP. If, however, two adjacent particles are of the same color but carry different charges, they can transform into of oppositely colored particles with different charges. Figure 1 provides a complete schematic of the transition rates.
@@
Figure 1: Transition rates of the particles in the model.
@@ 2.2.1 $R$-matrix
The $R$-matrix (2.36) can be obtained by doing a gauge transformation on the $R$ matrix in Ref. [14]. Let’s denote the $R$ matrix in Ref. [14] as $\tilde{R}$, then we have $R=(F\otimes F)\,\tilde{R}\,(F^{-1}\otimes F^{-1})$ with $F={\rm diag}(1,1,-1,1)$.
@@ 2.3.1 Decomposition of the transfer matrix
The first order derivative of $\ln t_{\sigma,\tau}(u)$ gives the Markov transition matrix of SSEP with periodic boundary (or the Hamiltonian of a periodic XXX chain)
@@ 2.3.2 Exact solutions
The steady states of the model read
@@
Figure 5: Dependence of the expansion coefficient on time $t$ in the regime $t\ll 1$. The initial state is $|-2\rangle\otimes|-1\rangle\otimes|+1\rangle$ with $N=3$. Left panel: Time evolution of the coefficient $c_{-2,-2,+2}$; the dashed line corresponds to $y=t$. Right panel: Time evolution of the coefficient $c_{-1,+1,-2}$; the dashed line corresponds to $y=t^{2}$.
3 Integrable stochastic process with twisted boundary condition
@@ 3.3.1 Decomposition of the transfer matrix
The transition matrix in Eq. (3.2) is the direct sum of $\overline{\mathcal{M}}_{\sigma}$ and $\overline{\mathcal{M}}_{\tau}$
@@
which implies that the model can be decomposed into two independent SSEPs with twisted boundaries.
@@ 3.3.2 Exact solutions
$\langle\hat{n}_{j_{1},s_{1}}\hat{n}_{j_{2},s_{2}}\cdots\hat{n}_{j_{n},s_{n}}\rangle$ $=\frac{\langle\Psi|\hat{n}_{j_{1},s_{1}}\hat{n}_{j_{2},s_{2}}\cdots\hat{n}_{j_{n},s_{n}}|\Psi\rangle}{\langle\Psi|\Psi\rangle}$
$=\frac{\sum_{k^{\prime},k=1}^{4}c_{k^{\prime}}{c_{k}}\langle\Psi_{k^{\prime}}|\hat{n}_{j_{1},s_{1}}\hat{n}_{j_{2},s_{2}}\cdots\hat{n}_{j_{n},s_{n}}|\Psi_{k}\rangle}{\sum_{k^{\prime},k=1}^{4}c_{k^{\prime}}{c_{k}}\langle\Psi_{k^{\prime}}|\Psi_{k}\rangle}$
$=\frac{\sum_{k=1}^{4}{c^{2}_{k}}\langle\Psi_{k}|\hat{n}_{j_{1},s_{1}}\hat{n}_{j_{2},s_{2}}\cdots\hat{n}_{j_{n},s_{n}}|\Psi_{k}\rangle}{\sum_{k=1}^{4}{c^{2}_{k}}\langle\Psi_{k}|\Psi_{k}\rangle}=\frac{1}{4^{n}},$ (3.63)
where $j_{l}=1,\ldots,N,\,\,j_{l}\neq j_{l^{\prime}},$ and $s_{l}=\pm 2,\pm 1$.
@@
4 Integrable stochastic process with open boundary conditions

### v2

@@ Abstract
We introduce an integrable stochastic process associated with the $D_{2}$ quantum group, which can be decomposed into two symmetric simple exclusion processes. We establish the integrability of the model under three types of boundary conditions (periodic, twisted, and open boundaries), and present its exact solution, including the spectrum, eigenstates, and some observables. This integrable model can be generalized to the asymmetric case, decomposing into two asymmetric simple exclusion processes, and its exact solutions are also studied.
@@ 1 Introduction
In this work, we construct a novel integrable stochastic process associated with the $D_{2}$ spin chain [14, 15], which represents a prototypical quantum integrable system not based on an $A$-type Lie algebra, and has many applications in the high energy, topological and mathematical physics. Our model incorporates four types of particles distinguished by their “charge” and “color”. Owing to the Lie algebra isomorphism $SO(4)\cong SU(2)\oplus SU(2)$ (often denoted as $D_{2}\cong A_{1}\oplus A_{1}$ in Cartan classification), the system can be decomposed into two independent SSEPs (or equivalently, $SU(2)$ XXX spin chains). We demonstrate the integrability of the model under periodic, twisted, and open boundary conditions and derive its exact solution in the form of TTQQ relation, Bethe state and associated Bethe ansatz equations. Our model can be generalized to the asymmetric case, where the system can be decomposed into two ASEPs. In this scenario, the system may be associated with the $D_{2}^{(1)}$ spin chain [16, 17, 18]. The models introduced in this paper differ from existing multi-component stochastic processes with integrable boundary conditions, thereby enriching the family of exactly solvable stochastic systems.
The paper is structured as follows: In Section 2, Section 3, and Section 4, we study the stochastic process with periodic, twisted, and open boundary conditions, respectively. For each case, we demonstrate the integrability of the model and derive the corresponding exact solutions. We generalize the models to the asymmetric case and present their exact solutions in Section 5. The paper concludes with a summary and outlook in Section 6.
2 Integrable symmetric stochastic process with period boundary conditions
@@ 2.1 Transition matrix
When two particles of different colors occupy adjacent sites, they exchange positions with a symmetric transition rate, analogous to the dynamics in SSEP (see Figure 1(a)). When particles at adjacent lattice sites have the same color but different charges, they can transform into particles whose color is different from that of the original particles, and the charges of these new particles are distinct from each other (see Figure 1(b)).
@@
Figure 1: Transition rates of the particles in the model. Here $i,j,k=\pm 1,\pm 2$ denote the “charge” values of the particles, and different colors in the figure are used to label the particle “color”.
@@ 2.2.1 $R$-matrix
The $R$-matrix (2.36) can be obtained by applying a gauge transformation to the $R$ matrix associated with the $D_{2}$ algebra in Ref. [14]. Let’s denote the $R$ matrix in Ref. [14] as $\tilde{R}$, then we have $R=(F\otimes F)\,\tilde{R}\,(F^{-1}\otimes F^{-1})$ with $F={\rm diag}(1,1,-1,1)$.
@@ 2.3.1 Decomposition of the transfer matrix
Remark.
The decomposition of the $R$-matrix in Eq. (2.43) requires further clarification. One can expand the matrix $\mathcal{R}^{\sigma}_{\bar{1},\bar{2}}(u)$ in the space $V_{\bar{1}}\otimes V_{\bar{2}}$ as follows
$\mathcal{R}^{\sigma}(u)=\sum_{k=1}^{4}\gamma_{k}(u)z(k)\otimes z(k),$ (2.49)
$\gamma_{1}(u)=\gamma_{2}(u)=\gamma_{3}(u)=\tfrac{1}{2},\quad\gamma_{4}(u)=u+\tfrac{1}{2},$
$z(1)=\sigma_{x},\quad z(2)=\sigma_{y},\quad z(3)=\sigma_{z},\quad z(4)=\mathbb{I}.$
Consequently, from Eq. (2.43), the $R$-matrix $R_{1,2}(u)$ in (2.36) (acting on $V_{1}\otimes V_{2}$) can be written in the form
$R(u)=\sum_{k,k^{\prime}=1}^{4}\gamma_{k}(u)\gamma_{k^{\prime}}(u)\underbrace{\left(z(k)\otimes z(k^{\prime})\right)}_{V_{1}}\otimes\underbrace{\left(z(k)\otimes z(k^{\prime})\right)}_{V_{2}}.$ (2.50)
@@
The first order derivative of $\ln t_{\sigma,\tau}(u)$ gives the Markov transition matrix of SSEP with periodic boundary (or equivalently, the Hamiltonian of a periodic XXX spin chain)
@@ 2.3.2 Exact solutions
In the limit $u\to\infty$, the rescaled operator $\mathcal{B}_{s}(u)/u^{N-1}$ converges to the total spin‑lowering operator
$\lim_{u\to\infty}\frac{\mathcal{B}_{s}(u)}{u^{N-1}}=\sum_{n=1}^{N}\sigma_{n}^{-},\quad s=\sigma,\tau.$ (2.69)
Consequently, the steady states of the model read
@@
Figure 5: Dependence of the expansion coefficient $c_{j_{1},\ldots,j_{N}}$ on time $t$ in the regime $t\ll 1$. The initial state is $|-2\rangle\otimes|-1\rangle\otimes|+1\rangle$ with $N=3$. Left panel: Time evolution of the coefficient $c_{-2,-2,+2}$; the dashed line corresponds to $y=t$. Right panel: Time evolution of the coefficient $c_{-1,+1,-2}$; the dashed line corresponds to $y=t^{2}$.
3 Integrable symmetric stochastic process with twisted boundary condition
@@ 3.3.1 Decomposition of the transfer matrix
The transition matrix $\mathcal$$\overline${M}} is the direct sum of $\overline{\mathcal{M}}_{\sigma}$ and $\overline{\mathcal{M}}_{\tau}$
@@ 3.3.2 Exact solutions
$\langle\hat{n}_{j_{1},\alpha_{1}}\hat{n}_{j_{2},\alpha_{2}}\cdots\hat{n}_{j_{n},\alpha_{n}}\rangle$ $=\frac{\langle\Psi|\hat{n}_{j_{1},\alpha_{1}}\hat{n}_{j_{2},\alpha_{2}}\cdots\hat{n}_{j_{n},\alpha_{n}}|\Psi\rangle}{\langle\Psi|\Psi\rangle}$
$=\frac{\sum_{k^{\prime},k=1}^{4}c_{k^{\prime}}{c_{k}}\langle\Psi_{k^{\prime}}|\hat{n}_{j_{1},\alpha_{1}}\hat{n}_{j_{2},\alpha_{2}}\cdots\hat{n}_{j_{n},\alpha_{n}}|\Psi_{k}\rangle}{\sum_{k^{\prime},k=1}^{4}c_{k^{\prime}}{c_{k}}\langle\Psi_{k^{\prime}}|\Psi_{k}\rangle}$
$=\frac{\sum_{k=1}^{4}{c^{2}_{k}}\langle\Psi_{k}|\hat{n}_{j_{1},\alpha_{1}}\hat{n}_{j_{2},\alpha_{2}}\cdots\hat{n}_{j_{n},\alpha_{n}}|\Psi_{k}\rangle}{\sum_{k=1}^{4}{c^{2}_{k}}\langle\Psi_{k}|\Psi_{k}\rangle}=\frac{1}{4^{n}},$ (3.63)
where $j_{l}=1,\ldots,N,\,\,j_{l}\neq j_{l^{\prime}},$ and $\alpha_{l}=\pm 2,\pm 1$.
@@
4 Integrable symmetric stochastic process with open boundary conditions